OpenStudy (anonymous):

A step-down transformer has a turn ratio of 6:1. A 240 volt alternating current supply at a frequency of 50 Hz is connected to the primary coil. What is the frequency of the AC output?

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

sorry about the double post but no replies yet :(

OpenStudy (anonymous):

Transformer does not (CANNOT) change frequency. Therefore, frequency will remain same i.e 50 Hz.(same as input). Here are some formulae for the transformer: N1/N2 = V1/V2 = I2/I1, where N => no. of turns V => Voltage I => Current 1 => for primary i.e input 2 => for secondary i.e. output. No relation for frequency. The above explanation is considering an ideal transformer. In real life, there may be very very little change in frequencies DUE TO LOSSES.

OpenStudy (anonymous):

thanks XD